Suppose that you want to create a portfolio that consists of a corporate bond​ fund, X, and a common stock​ fund, Y. For a​ $1,000 investment, the expected return for X is $ $77 and the expected return for Y is $ 91. The variance for X is 1,375 and the variance for Y is 13,850. The covariance of X and Y is 4,624. Complete parts​ (a) through​ (d).

a. Compute the portfolio expected return and portfolio risk if you put
$ 300$300
in the corporate bond fund and
$ 700$700
in the common stock fund.
Answer #1

Solution :-

Given data:

Total investment = $1,000

Expected return for X = $77

Expected return for Y = $ 91

Variance for X, Var_x =1,375

Variance for Y, Var_{y} = 13,850

Co-variance of X and Y, C.V_{xy} = 4,624

\RightarrowHere we have to calculate the portfolio expected return and portfolio risk.

Given as, X = $300 and Y = $700

\Rightarrow Now consider the tabular form,

Security X Y
Return

= $77/ $1,000

= 0.077

= 7.7%

= $ 91 / $1,000

= 0.091

= 9.1%

Risk (S.D=\sqrt{Var})

= \sqrt{1,375 }

= 37.0809

=  \sqrt{13,850}

= 117.6860

C.V_{xy}

4,624

\Rightarrow Now we have to calculate the Coefficient of correlation.

Here we know the formula, i.e

Coefficient of correlation = \frac{ C.V_{xy }}{ SD of X * SD of Y }

By substituting values we get,

Coefficient of correlation = \frac{ 4624 }{ 37.0809 * 117.6860 }

Coefficient of correlation = 1.0596

\Rightarrow Now consider the tabular form,

Security Weights (w) Return (R) W * R
X

= 300 / 1000

= 0.30

= 0.077

= 0.30 * 0.077

= 0.0231

Y

= 700 / 1000

= 0.70

= 0.091

= 0.70 *  0.091

= 0.0637

Return

= 0.0231 + 0.0637

= 0.0868

= 0.0868 * 100

= 8.68%

\therefore Expected return = 8.68%

Now consider the Variance of the portfolio, to find out the risk.

Variance of the portfolio = \left (W_{X} \right )^{2}* Var_x + \left (W_{Y} \right )^2* Var_y + 2* SD_x * SD_y * W_x * W_y * \left (Coefficient Of Correlation \right )xy

By substituting values we get,

Variance of the portfolio =  (0.30 )^{2}*1375 +( 0.70 )^{2}*13850 + 2* 37.0809 * 117.6860* 0.30 *0.70* 1.0596

Variance of the portfolio = 123.75 +6786.5+1942.07639

Variance of the portfolio = 8852.32

Then, Risk = \sqrt{Variance }

Risk = \sqrt{8852.32}

Risk = 94.08680

\therefore Risk = 94.08680